OpenStudy (haldko):

I just need someone to check my answer! I'll fan and medal Find the derivative of f(x) = 5 divided by x at x = -1.

OpenStudy (haldko):

I said either 5 or -1

OpenStudy (haldko):

@mathmate @FaiqRaees @Hero

OpenStudy (faiqraees):

=d/dx 5/x when x =-1 =5 (d/dx 1/x) Can you figure out the derivative of 1/x

OpenStudy (haldko):

I'm not sure @FaiqRaees

OpenStudy (faiqraees):

Are you aware of the formula ? \[\large\rm \frac{d}{dx}x^n =nx^{n-1} \]

OpenStudy (haldko):

No, I'm really new at these

OpenStudy (haldko):

@FaiqRaees

OpenStudy (haldko):

Pleasee help

OpenStudy (faiqraees):

Can you understand the notation above?